The Distributive Property (Fractions)
1/8 (x-16)= 20

Answer by checkley71(8403)

1/8(X-16)=20
X/8-2=20
X/8=22
X=22*8
X=176
PROOF
1/8(176-16)=20
1/8*160=20
20=20